Triple Strike, owned and managed by Mr. Brian, showed owner’s equity as of January 1, 2019 as P120,000. It is estimated that at December 31, 2019, total assets will be P400,000 and total liabilities P200,000. How much should be the 2019 projected profit or loss given the following additional information? Each case is to be considered as independent from the others:

     a. Brian did not make any additional investment or withdrawal.

     b. Brian made a cash withdrawal of P50,000 but no additional investment.

     c.  Brian made an additional investment of P125,000 but no cash withdrawal.

     d. Brian made an additional investment of P125,000 and a cash withdrawal of P50,000.



1
Expert's answer
2021-02-22T13:59:52-0500

Net Amount of Capital Before Profit  = Brian, Capital Jan. 1 + Investment - Drawings 

Brian, Capital Dec. 31 = total assets - total liabilities = P400,000 - P200,000 = P200,000

Profit or Less Loss = Brian, Capital Dec. 31 - Net Amount of Capital Before Profit

a. Profit or Less Loss = 200,000 - (120,000 + 0 - 0) = P80,000

b. Profit or Less Loss = 200,000 - (120,000 + 0 - 50,000) = P130,000

c. Profit or Less Loss = 200,000 - (120,000 + 125,000 - 0) = -P45,000

d. Profit or Less Loss = 200,000 - (120,000 + 125,000 - 50,000) = P5,000